OpenStudy (anonymous):

The points (-9, r) and (-5, 13) lie on a line with the slope 2. Find the missing coordinate r. .

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{13-r}{-5--9}=slope\] \[\LARGE \frac{13-r=?}{-5+9=?}=slope\]

OpenStudy (anonymous):

13-r/4

pooja195 (pooja195):

13-r/4=slope What should we subtract from 13 so that the slope will be 2?

OpenStudy (anonymous):

4

pooja195 (pooja195):

13-r=4 13-4=?

OpenStudy (anonymous):

9

pooja195 (pooja195):

Good so we get r=9 Because 13-9=4/2=2